• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 21:34:11 on September 01, 2009

    Entry number 287393

    keyword=Compton Electron detector status

    No signal seen since access. Not sure if the beam moved. Detector current seems fine. Turned off left side, it seems the part which had a power supply issue is very noisy. A few electrons seen on the analog board but less obvious than last time. Did a scan a 2 4 and 6 mm, did not see any obvious Compton events. I leave it at 6mm with photon trigger firmware.

    A copy of this log entry has been emailed to: rom@jlab.org



    Figure 1